# Large time behavior for the 3D Navier-Stokes with Navier boundary conditions

### v3

@@ 8. Comments and conclusions
Our main results are the existence of a Leray-type weak solution, with two versions of the corresponding energy inequality, and three long-time exponential decay estimates. The basic structure of the arguments are classical. Still, beyond closing a gap in the literature, the main point of this work is to account for the influence of the differential geometry of the boundary on this problem. This arises in several moments:
@@
Much of our analysis is inspired by the two-dimensional work done by Clopeau et al in [CMR]. Notably, the formulation of the Navier boundary conditions which makes explicit the influence of the geometry of the boundary is already present: see [CMR, Lemma 2.1]. Furthermore, their work depends on the construction of an appropriate Galerkin basis, which inspired the corresponding construction presented here.
@@ Acknowledgments
Part of this work was prepared while Kelliher was participating in a program hosted by the Simons Laufer Mathematical Sciences Research Institute in Berkeley, California, in Spring 2021 and again in Summer 2023, supported by the National Science Foundation under Grant No. DMS-1928930. The second author is partially supported by the French National Research Agency in the framework of the project “BOURGEONS” (ANR-23-CE40-0014-01). The third and fourth authors gratefully acknowledge the hospitality of the Department of Mathematics at the University of California, Riverside, and of the Institut Fourier at the Université de Grenoble, where part of this research was done. MCLF was partially supported by CNPq, through grant # 304990/2022-1, and FAPERJ, through grant # E-26/201.209/2021. HJNL acknowledges the support of CNPq, through grant # 305309/2022-6, and of FAPERJ, through grant # E-26/201.027/2022. The work of E.S.T. was partially supported by the DFG Research Unit FOR 5528 on Geophysical Flows. In addition, MCLF and HJNL would like to thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the program “Mathematical aspects of turbulence: where do we stand”, when part of the work on this paper was undertaken. This work was supported in part by: EPSRC Grant Number EP/R014604/1.

### v4

@@ 8. Comments and conclusions
Our main results are the existence of a Leray-type weak solution, with two versions of the corresponding energy inequality, and three long-time exponential decay estimates. The basic structure of the arguments are classical. Still, aside from providing a comprehensive account of existence and large-time behavior of weak solutions for Navier-Stokes with Navier boundary conditions, the main point of this work is to account for the influence of the differential geometry of the boundary on this problem. This arises in several moments:
@@
Much of our analysis is inspired by the two-dimensional work done by Clopeau et al in [CMR]. Notably, the formulation of the Navier boundary conditions which makes explicit the influence of the geometry of the boundary is already present: see [CMR, Lemma 2.1]. Furthermore, their work depends on the construction of an appropriate Galerkin basis, which inspired the corresponding construction presented here. We note that, in [FG2022DCDS], an outline of the Galerkin approximation was obtained in the case $\alpha=0$ and for three-dimensional Lipschitz domains.
@@ Acknowledgments
Part of this work was prepared while Kelliher was participating in a program hosted by the Simons Laufer Mathematical Sciences Research Institute in Berkeley, California, in Spring 2021 and again in Summer 2023, supported by the National Science Foundation under Grant No. DMS-1928930. The second author is partially supported by the French National Research Agency in the framework of the project “BOURGEONS” (ANR-23-CE40-0014-01) and “ComplexFlows” of the PEPR MathsViVEs (ANR-23-EXMA-0004). The third and fourth authors gratefully acknowledge the hospitality of the Department of Mathematics at the University of California, Riverside, and of the Institut Fourier at the Université de Grenoble, where part of this research was done. MCLF was partially supported by CNPq, through grant # 304990/2022-1, and FAPERJ, through grant # E-26/201.209/2021. HJNL acknowledges the support of CNPq, through grant # 305309/2022-6, and of FAPERJ, through grant # E-26/201.027/2022. The work of E.S.T. was partially supported by the DFG Research Unit FOR 5528 on Geophysical Flows. In addition, MCLF and HJNL would like to thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the program “Mathematical aspects of turbulence: where do we stand”, when part of the work on this paper was undertaken. This work was supported in part by: EPSRC Grant Number EP/R014604/1.